OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. cos x - sin 2x - 1

OpenStudy (anonymous):

\[\cos(x)-\sin(2x)-1\] or \[\cos(x)-\sin^2(x)-1\]

OpenStudy (anonymous):

@satellite73 can try my question pls

OpenStudy (anonymous):

It's cos(x)-sin^2(x)-1

OpenStudy (anonymous):

ok then maybe we can use \[\sin^2(x)+\cos^2(x)=1\] somewhere

OpenStudy (anonymous):

replace \(\sin^2(x)\) by \(1-\cos^2(x)\) in your expression you get \[\cos(x)-(1-\cos^2(x))-1\] then some algebra to fix it up a bit